32.46 ml of nitrogen measured at STP is relesed when 0.1g of a primary amine is acted upon by excess of nitrous acid..       R-NH2+HNOgives R-OH+N2+H2O.calculate molar mass of amine..



Grade:11

1 Answers

abhishek upadhyay
12 years ago

let its molar mass be m then by stiochiometry

0.1/m = 32.46/(1000*22.4)   or m = 69g